OpenStudy (anonymous):

(-8)^2/3 ? I'm the worst with fractions...

OpenStudy (anonymous):

What is negative 8 to the power of 2/3?

OpenStudy (anonymous):

That's what it's asking.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

-21.3333333

OpenStudy (jdoe0001):

\(\large { \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \\ \quad \\ \quad \\ (-8)^{\frac{2}{3}}\implies \sqrt[{\color{red}{ 3}}]{(-8)^{\color{blue}{ 2}}} } \)

OpenStudy (anonymous):

That's what I thought but the only answers I have to choose from is: 4, -4, -2 & -1/2

OpenStudy (freckles):

what is the square of -8?

OpenStudy (freckles):

or in other words what is -8(-8)?

OpenStudy (anonymous):

\[-8^{2/3}\]

OpenStudy (freckles):

\[\sqrt[3]{(-8)(-8)}=\sqrt[3]{64}=\sqrt[3]{4 \cdot 4 \cdot 4 }=?\] i will let you finish this

OpenStudy (anonymous):

Thank you!!!!